Two particles, with charges of 20.0 nC and 40.0 nC are placed at the points with coordinates (0, 4.00 cm) and (3.00 cm, 0). A particle with charge 10.0 nC is located at the origin. (a) Find the electric potential energy of the configuration of the three fixed charges.
1
Expert's answer
2020-10-23T07:16:39-0400


Let "20.0\ nC" charge be "q_1", "40.0\ nC" charge be "q_2" and "10.0\ nC" charge be "q_3". Let, also, the distance between the charges "q_1"and "q_2" be "r_{12}", the distance between the charges "q_2"and "q_3" be "r_{23}" and the distance between the charges "q_1"and "q_3" be "r_{13}".

Let's write the formula for the electric potential energy due to charges "q_1"and "q_2":


"U_1=k\dfrac{q_1q_2}{r_{12}}=k\dfrac{q_1q_2}{\sqrt{(r_{13}^2+r_{23}^2)}},"

here, "k=8.99\cdot 10^9 \ \dfrac{Nm^2}{C^2}" is Coulomb constant, "r_{13}=4\cdot 10^{-2}\ m" is the distance between the charges "q_1"and "q_3", "r_{23}=3\cdot 10^{-2}\ m" is the distance between the charges "q_2"and "q_3".

Then, we can calculate "U_1":

"U_1=8.99\cdot 10^9 \ \dfrac{Nm^2}{C^2}\cdot \dfrac{20.0\cdot 10^{-9}\ C \cdot 40.0\cdot 10^{-9}\ C}{\sqrt{((4\cdot 10^{-2}\ m)^2+(3\cdot 10^{-2}\ m)^2)}}=1.44\cdot 10^{-4}\ J."

Similarly, we can write the formula for the electric potential energy due to charges "q_2"and "q_3":


"U_2=k\dfrac{q_2q_3}{r_{23}}."


Let's calculate "U_2":

"U_2=8.99\cdot 10^9 \ \dfrac{Nm^2}{C^2}\cdot\dfrac{40.0\cdot 10^{-9}\ C \cdot 10.0\cdot 10^{-9}\ C}{3\cdot 10^{-2}\ m}=1.2\cdot 10^{-4}\ J."

Finally, we can write the formula for the electric potential energy due to charges "q_1"and "q_3":


"U_3=k\dfrac{q_1q_3}{r_{13}}."


Let's calculate "U_3":

"U_3=8.99\cdot 10^9 \ \dfrac{Nm^2}{C^2}\cdot\dfrac{20.0\cdot 10^{-9}\ C \cdot 10.0\cdot 10^{-9}\ C}{4\cdot 10^{-2}\ m}=0.45\cdot 10^{-4}\ J."

Then, the net electric potential energy of the configuration of the three fixed charges is the sum of "U_1", "U_2" and "U_3":


"U_{net}=U_1+U_2+U_3,"

"U_{net}=1.44\cdot 10^{-4}\ J+1.2\cdot 10^{-4}\ J+0.45\cdot 10^{-4}\ J=3.1\cdot 10^{-4}\ J."

Answer:

"U_{net}=3.1\cdot 10^{-4}\ J."